OpenStudy (anonymous):

Find the derivative of the function y=(4x+3)^4 (x+1)^-3

OpenStudy (jdoe0001):

I'd think the product rule will do

OpenStudy (jdoe0001):

\(\bf [f(x)\cdot g(x)]' \implies f(x)'\cdot g(x) + g(x)'\cdot f(x)\)